### v1

# Quantifying the Impact of Incompleteness on Identifying and Interpreting Galaxy Protocluster Populations with the TNG-Cluster Simulation

Abstract
We use the TNG-Cluster simulation to examine how stellar mass and star formation rate (SFR) incompleteness affect the identification of density peaks within galaxy protoclusters at different redshifts. We focus on 352 protoclusters, defined as the set of galaxies that will reside within the virialized region of a $z=0$ cluster halo with total mass $\sim 10^{14.3-15.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, and consider only galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ (our baseline) at any redshift. We find that ${M}_{\star}$-limited (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) and SFR-limited ($\mathrm{SFR}>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) subpopulations only recover the true highest galaxy density peak in $\lesssim 40\%$ of cases within an accuracy of $1.0$ pMpc ($\sim 2-2.5\arcmin$) at $z>2$. We find that at $z>2$ the highest galaxy density peaks do not generally coincide with the highest dark matter or stellar mass density peaks, and this separation is larger than 0.5 pMpc in $\sim 80\%$ of cases. Consequently, the region surrounding the true highest galaxy density peaks are not generally sites of enhanced star formation or accelerated mass growth relative to the remainder of the protocluster. Lastly, we find that the $\sim 4\arcmin$ apertures typically used to define spectroscopically-confirmed protoclusters are generally much smaller than the $8\arcmin$ apertures needed to study the progenitors of the most massive galaxy clusters at $z>2$.
High-redshift galaxy clusters (2007), Galaxy formation (595), Galaxy evolution (594), Galaxy environments (2029), Large-scale structure of the universe (902)
software: astropy (Astropy Collaboration et al., 2013b, 2018b)
1 Introduction
Galaxy clusters are the most massive gravitationally bound structures in the universe, with masses surpassing one hundred trillion Suns and harboring hundreds to thousands of galaxies. In the standard $\Lambda$CDM model of cosmology these rare cosmic behemoths form over billions of years, emerging from tiny fluctuations in the initial cold dark matter density field and growing hierarchically in an expanding universe through mergers and accretion (Peebles & Yu, 1970; Press & Schechter, 1974; Bond et al., 1996; Schneider, 2015). As they assemble galaxy clusters eventually collapse and reach a dynamically relaxed (or “virialized”) state, characterized by the presence of a superheated ($\sim 10^{8}$ K) plasma known as the intracluster medium (ICM) (Sarazin, 1986; Briel et al., 1992; Elbaz et al., 1995) and by the presence of a population of massive red elliptical galaxies that constitute the “red sequence” (Bower et al., 1992; Gladders & Yee, 2000).
In addition to being extreme and rare cosmic structures, galaxy clusters serve as valuable laboratories for studying the impact of dense environments on galaxy evolution (e.g., Oemler, 1974; Dressler, 1980; Moore et al., 1996; Kauffmann et al., 2004; Cooper et al., 2006; Peng et al., 2010; Wetzel et al., 2013; Lemaux et al., 2017; Baxter et al., 2022; Kukstas et al., 2023; Baxter et al., 2023; Taamoli et al., 2024), investigating the nature and distribution of dark matter (e.g., Clowe et al., 2006; Bradač et al., 2006, 2008; Umetsu et al., 2018; Wittman et al., 2023), and constraining cosmological parameters (e.g., Birkinshaw & Hughes, 1994; Carlberg et al., 1996; Eke et al., 1996; Schuecker et al., 2003; Benson et al., 2013; Hung et al., 2021).
While the late stages of cluster assembly have been extensively studied through observations of virialized (or near-virialized) clusters to $z\lesssim 1.5$ (e.g., Ebeling et al., 2001; Lubin et al., 2009; Wilson et al., 2009; Muzzin et al., 2012; Reichardt et al., 2013; Gonzalez et al., 2019; Golden-Marx et al., 2019; Balogh et al., 2021; Biviano et al., 2021), the pre-virialized or “protocluster” stage remains less well understood. This is partly because observations of virialized clusters provide limited insight into their early formation history. Transformative events such as mergers and dynamical relaxation significantly reshape clusters and their galaxy populations, often erasing evidence of their initial conditions — though dynamical analysis can still provide clues about their assembly history (Biviano & Katgert, 2004). This challenge is further compounded by the difficulty of identifying and characterizing galaxy protoclusters, which lack clear markers characteristic of virialized galaxy clusters at $z\lesssim 2$, such as an established ICM or a prominent red sequence. For example, the ICM is crucial for identifying and characterizing galaxy clusters, as it emits X-rays via thermal bremsstrahlung (free-free emission) and produces a distinctive decrement in the cosmic microwave background (CMB) at submillimeter wavelengths through the thermal Sunyaev-Zel’dovich (SZ) effect (Sunyaev & Zeldovich, 1970), caused by high-energy ICM electrons upscattering CMB photons. These signatures of ICM physics are essential for detecting galaxy clusters in X-ray (e.g., Takey et al., 2016; Koulouridis et al., 2021; Xu et al., 2022) and SZ (e.g., Bleem et al., 2015, 2024; Klein et al., 2024) surveys.
Without these distinctive features, galaxy protoclusters are predominantly identified via high-redshift ($z>2$) galaxy overdensities. These overdensities have been traced using a wide range of galaxy populations including Lyman-break galaxies (LBGs) (e.g., Steidel et al., 1998; Brinch et al., 2024; Toshikawa et al., 2025a), H-alpha emitters (HAEs) (e.g., Shi et al., 2021a; Pérez-Martínez et al., 2023), and Lyman-alpha emitters (LAEs) (e.g., Venemans et al., 2007; Overzier et al., 2008; Harikane et al., 2019), which are representative of normal star-forming galaxies at early times. Other tracers include highly luminous galaxy populations, such as dusty star-forming galaxies (DSFGs) (e.g., Oteo et al., 2018; Gómez-Guijarro et al., 2019; Long et al., 2020) and submillimeter galaxies (SMGs) (e.g., Casey et al., 2015; Jones et al., 2017; Lacaille et al., 2019; Calvi et al., 2023). Protoclusters have also been identified through overdensities in photometric and spectroscopic surveys (e.g., Lemaux et al., 2014; Franck & McGaugh, 2016a, b; Toshikawa et al., 2020; Hung et al., 2025; Toni et al., 2025), as well as via biased tracers of high-redshift overdensities, such as radio galaxies (e.g., Hatch et al., 2011; Hayashi et al., 2012; Cooke et al., 2014; Shen et al., 2021), quasars (e.g., Djorgovski et al., 2003; Morselli et al., 2014; García-Vergara et al., 2017), and ultra-massive galaxies (UMGs) (McConachie et al., 2022, 2025). Beyond high-$z$ galaxy overdensities, protoclusters have also been identified through large-scale intergalactic medium (IGM) overdensities, traced by Lyman-alpha absorption from UV light emitted by background quasars and LBGs interacting with intervening neutral hydrogen (Lee et al., 2014, 2016; Cai et al., 2017; Lee et al., 2018; Newman et al., 2020).
The current sample of spectroscopically-confirmed galaxy protoclusters remains generally sparse and heterogeneously selected. While some consistent trends have emerged – such as protocluster environments, relative to coeval field populations, being regions of accelerated galaxy evolution (Steidel et al., 2005; Hatch et al., 2011; Koyama et al., 2013; Cooke et al., 2014; Shimakawa et al., 2018a; Forrest et al., 2024; Helton et al., 2024a), enhanced star formation rates (Dannerbauer et al., 2014; Hayashi et al., 2016; Miller et al., 2018), enhanced AGN activity (Digby-North et al., 2010; Lemaux et al., 2014; Tozzi et al., 2022; Shah et al., 2024), enhanced merger rates (Hine et al., 2016; Liu et al., 2023; Giddings et al., 2025), and sites of a reversal in the star formation rate–density relation (Tran et al., 2010; Popesso et al., 2011; Lemaux et al., 2022) – other trends remain contentious or poorly understood. For instance, it is uncertain whether protoclusters are metal-enriched, metal-deficient, or neither (Kulas et al., 2013; Alcorn et al., 2019; Sattari et al., 2021), and the extent to which protocluster environments suppress or “quench” star formation is still unclear (e.g., Kubo et al., 2021; Shi et al., 2021b; Alberts & Noble, 2022; McConachie et al., 2022; Tanaka et al., 2024; Edward et al., 2024; Espinoza Ortiz et al., 2024; Naufal et al., 2024; McConachie et al., 2025).
More broadly, our understanding of protoclusters remains largely phenomenological, with investigations primarily limited to characterizing differences between protoclusters and coeval field populations, while the physics driving this differential evolution remains elusive. This limitation arises not only from small and heterogeneous sample sizes, which hinder population-level analyses, but also from fundamental challenges in identifying protoclusters, constructing unbiased samples, and acquiring sufficient spectroscopic data over large areas to map their full extent and assemble a tracer population that is representative of the underlying galaxy distribution. However, efforts to build large, uniformly selected, and spectroscopically-confirmed protocluster samples are underway (Hung et al., 2025). These efforts, combined with ever-growing spectroscopic catalogs in legacy fields Khostovan et al. (2025) and ongoing and upcoming surveys with next-generation observatories – including JWST (Li et al., 2024), Euclid (Euclid Collaboration et al., 2025a), LSST (Ivezić et al., 2019; Gully et al., 2024), and Roman (Rudnick et al., 2023) – will play a significant role in advancing this field.
Cosmological simulations of galaxy formation serve as powerful tools for interpreting existing protocluster observations, making testable predictions for future surveys, and guiding their development. This is largely due to simulations providing a self-consistent framework that bridges the gap between high-redshift galaxy overdensities and virialized clusters at $z=0$. Simulations generally use this connection to define protoclusters as collections of galaxies that will eventually reside in virialized clusters at $z=0$. Following this definition, Chiang et al. (2013) and Muldrew et al. (2015) used the Millennium Run (MR) dark matter simulation (Springel et al., 2005) with semi-analytic models of galaxy evolution to study protocluster evolution. Chiang et al. (2013) analyzed the mass and size evolution of $\sim 3000$ protoclusters, predicting that overdensities at $z<5$ correlate with the halo mass of the cluster at $z=0$. Muldrew et al. (2015) further demonstrated that the evolutionary state of a protocluster can be inferred from the mass ratio of its two most massive galaxies. Expanding on this, Chiang et al. (2017) predicted that protoclusters contribute significantly to cosmic star formation rate density at high redshifts – about $20\%$ at $z=2$ and up to $50\%$ at $z=10$ – a prediction that has been supported observationally by Staab et al. (2024), based on an investigation of a single protostructure at $z\sim 4.5$. Similarly, Popescu et al. (2023) found comparable results for a stacked sample of Planck- identified protocluster candidates (Planck Collaboration et al., 2015) at $z=2-4$, further supporting these predictions.
Lim et al. (2021) used hydrodynamical simulations and empirical models to show that simulated protoclusters at $z>2$ underestimate observed star formation rates by at least a factor of three, attributing this to limitations in numerical resolution. However, Gouin et al. (2022), using the TNG300 simulation from the IllustrisTNG project (Nelson et al., 2019a), found that simulated protoclusters reproduce the star formation rates, stellar masses, and galaxy richness observed in Planck-selected high-redshift protoclusters. Lim et al. (2024) used the FLAMINGOS simulation suite (Schaye et al., 2023) to investigate how variations in aperture definitions affect mass estimates of protoclusters, finding that common observational aperture choices can introduce biases of up to an order of magnitude in total mass (baryonic and dark matter) estimates at $z\lesssim 4$. Using the DIANOGA zoom-in cosmological hydrodynamical simulations (Bonafede et al., 2011; Rasia et al., 2015), Esposito et al. (2025) found enhanced star formation suppression in protocluster galaxies at $z=2.2$, particularly in the most massive halos.
Other studies have leveraged constrained and large-scale simulations to investigate analogs of observed protoclusters. Ata et al. (2022) used constrained cosmological simulations, designed to match the observed galaxy distribution in the COSMOS field (Scoville et al., 2007), to confirm that many of these systems will collapse into massive clusters by $z=0$. Remus et al. (2023) utilized the Magneticum Pathfinder simulations (Dolag et al., 2016) to analyze the evolutionary history of protoclusters similar to SPT2349-56 (Miller et al., 2018) at $z\sim 4$, concluding that properties such as virial mass, star formation rate (SFR), stellar mass, and galaxy richness do not strongly correlate with the final cluster mass at $z=0$, and that these protoclusters will not be among the top ten most massive clusters at $z=0.2$. Yajima et al. (2022) used the FOREVER22 simulation to study the formation of supermassive black holes (SMBHs) and bright SMGs in SSA22 protocluster analogs at $z\sim 3$, finding that SMBHs form in the most massive halos (${M}_{\rm{halo}}\sim 10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$) and grow rapidly until their stellar mass exceeds ${M}_{\star}\gtrsim 10^{10}\leavevmode\nobreak\ {\rm M}_{\odot}$, after which feedback suppresses further growth, while dusty starburst galaxies emerge in massive halos (${M}_{\rm{halo}}\gtrsim 10^{13}\leavevmode\nobreak\ {\rm M}_{\odot}$) in the protocluster core.
In this work we use the TNG-Cluster simulation Nelson et al. 2024, which offers a unique combination of a large sample of very massive galaxy clusters (352 with $M_{\rm{200}}>10^{14.3}\leavevmode\nobreak\ {\rm M}_{\odot}$ at $z=0$) and high baryonic mass resolution ($1.2\times 10^{7}\leavevmode\nobreak\ {\rm M}_{\odot}$), to explore potential biases in observed protocluster populations. Specifically, we investigate how often the angular sizes used to probe spectroscopically-confirmed protoclusters align with their theoretically expected spatial extents. We also evaluate how stellar mass and SFR incompleteness affect the identification of the highest density peaks and compare the average properties of protocluster galaxies within and beyond these peaks.
This paper is structured as follows. In §2 we describe the TNG-Cluster simulation, define our protocluster samples, and explain our methods for quantifying the radial extent of protoclusters and their density fields. In §3 we present a comparison of predicted and observed sizes of galaxy protoclusters, quantify the impact of stellar mass and SFR incompleteness on identifying the highest density peaks in protoclusters, and compare the properties of protocluster galaxies within and beyond the highest density peak. Finally, in §4 we discuss the implications of our results and provide considerations for observational studies of galaxy protoclusters. We summarize our findings in §5.
Throughout this study we adopt a Planck 2015 cosmology with $H_{0}=67.7\leavevmode\nobreak\ {\rm km}{\rm s}^{-1}{\rm Mpc}^{-1}$ and $\Omega_{m}=0.307$ (Planck Collaboration et al., 2016), express distances in physical units, use only high-resolution particles from the zoom-in regions of the TNG-Cluster simulation, and determine quantities related to the dynamically relaxed (or virialized) region of clusters at the radius where the average density is 200 times the critical density of the Universe ($\rho_{\mathrm{crit}}=3H^{2}/8\pi G$).
2 Data and Methods
2.1 TNG-Cluster Simulation
In this study, we use the TNG-Cluster111www.tng-project.org/cluster simulation (Nelson et al., 2024), an extension of the IllustrisTNG suite of cosmological magnetohydrodynamical simulations of galaxy formation (Weinberger et al., 2017; Pillepich et al., 2018) run with the moving-mesh code Arepo (Springel, 2010). TNG-Cluster builds on TNG50 (Pillepich et al., 2019; Nelson et al., 2019b), TNG100, and TNG300 (Nelson et al., 2018; Pillepich et al., 2018; Springel et al., 2018; Marinacci et al., 2018; Naiman et al., 2018), significantly increasing the statistical sampling of the high-mass end of the halo mass function at $z=0$. This improved sampling of massive clusters results from TNG-Cluster being constructed from hundreds of multi-mass “zoom” re-simulations of cluster halos drawn from 1 Gpc volume. TNG-Cluster includes 352 clusters, producing $30$ times more clusters with $M_{\rm{200}}^{z=0}>10^{15}\leavevmode\nobreak\ {\rm M}_{\odot}$ than TNG300 while maintaining a mean baryonic (gas and stars) mass resolution of $m_{\mathrm{baryon}}=1.2\times 10^{7}\leavevmode\nobreak\ {\rm M}_{\odot}$, a dark matter mass resolution of $m_{\mathrm{DM}}=6.1\times 10^{7}\leavevmode\nobreak\ {\rm M}_{\odot}$, and a gravitational softening length of $1.5$ kpc for stars and dark matter at $z=0$. Its data products include particle-level snapshots, as well as halo and subhalo catalogs stored across $100$ snapshots from $z=15$ to $z=0$. Consistent with previous TNG simulations, dark matter halos are identified with the Friends-of-Friends algorithm (Davis et al., 1985), while the SUBFIND algorithm (Springel et al., 2001) identifies gravitationally-bound subhalos, with galaxies defined as subhalos with non-zero stellar mass. The SubLink merger tree (Rodriguez-Gomez et al., 2015) is used to track their evolution across time.
Figure 1: Left: The dark matter halo mass assembly history of the 352 galaxy clusters from the TNG-Cluster simulation. The gray lines show the individual assembly histories. The purple, blue, and green lines represent the median assembly histories, binned according to the clusters’ final masses at $z=0$, denoted as $M_{\rm{200}}^{z=0}$. The colored bands indicate the $68\%$ quantiles of the binned distribution. Right: The redshift at which each cluster first crosses the canonical halo mass threshold of $M_{200}>10^{14}\leavevmode\nobreak\ M_{\odot}$, a common criterion for defining galaxy clusters.
TNG-Cluster follows the IllustrisTNG physical model, incorporating gas radiative processes, star formation in the dense interstellar medium, stellar population evolution and chemical enrichment, supernova-driven galactic-scale outflows, and the formation, merging, and growth of SMBHs, including dual-mode SMBH feedback (see Nelson et al. 2024 for details). Initial investigations with TNG-Cluster have leveraged these aspects of the simulation to directly compare with observed X-ray cavities in galaxy clusters (Prunier et al., 2025a, b), provide theoretical expectations for the temperature and metallically distribution of the ICM (Chatzigiannakis et al., 2025), characterize the population of cool-core to non-cool-core clusters (Lehle et al., 2024) and identify the mechanisms driving the transformation from one to the other (Lehle et al., 2025), quantify the abundance, spatial distribution, and origin of cool gas in clusters (Rohr et al., 2024; Staffehl et al., 2025), provide theoretical expectations for the kinematics of gas (Truong et al., 2024; Ayromlou et al., 2024), in addition to the X-ray emitting gas content of massive cluster members (Rohr et al., 2024) and radio relics (Lee et al., 2024).
Figure 2: The projected spatial distribution of six clusters from the TNG-Cluster simulation, selected to fall within the three halo mass bins defined in Fig. 1. Each row represents an individual cluster, while each column corresponds to a different redshift snapshot. Black circles denote galaxies from our baseline population – galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within the virial radius ($R_{\rm{200c}}$, red circle) by $z=0$ – whereas gray circles represent galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside between $R_{\rm{200c}}<R<10\leavevmode\nobreak\ R_{\rm{200c}}$ at $z=0$.
2.2 Simulated Cluster Sample
In this study we analyze all $352$ galaxy clusters from TNG-Cluster to investigate their assembly histories and the influence of local galaxy density in protocluster environments on the properties of their constituent galaxies. In the left panel of Fig. 1 we show the halo mass assembly histories of all clusters (gray lines). The solid colored lines indicate the median halo mass for clusters grouped by their halo mass at $z=0$ (i.e., $M_{\rm{200}}^{z=0}$), with the shaded bands representing the $68\%$ quantiles.
A striking feature of this sample is the diversity in cluster assembly histories – some clusters accumulate their mass rapidly at early times, while others grow more gradually, only exceeding $10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$ at later times. However, as shown in the right panel of Fig. 1, a clear trend emerges when comparing the redshift at which a cluster first reaches the characteristic cluster mass scale ($\geq 10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$) and its halo mass at $z=0$. This trend reflects the hierarchical nature of structure formation, where more massive halos at $z=0$ tend to form earlier (though this trend reverses if “formation” is instead defined using a fractional mass threshold; Wechsler et al., 2002; Nadler et al., 2023). It also motivates partitioning the protocluster sample based on $M_{\rm{200}}^{z=0}$ and defining $z\sim 2$ as the transition epoch when the first halos surpass the characteristic cluster halo mass scale, marking the transition from galaxy protoclusters to clusters. Notably, the most distant known cluster, CL J1001 (Wang et al., 2016, 2018), located at $z=2.51$, is consistent with the redshift at which the first clusters in the TNG-Cluster simulation cross the canonical halo mass threshold of $M_{\rm{200}}>10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$.
2.3 Simulated Protocluster Population
In the review paper by Overzier (2016), galaxy protoclusters are defined as overdensities of galaxies that will eventually collapse into a galaxy cluster - i.e., a dynamically relaxed structure more massive than $10^{14}\leavevmode\nobreak\ {\rm M}_{\odot}$ at $z=0$. We adopt this definition to identify the protocluster population analyzed in this study, with the caveat that our minimum mass threshold is $10^{14.3}\leavevmode\nobreak\ {\rm M}_{\odot}$. Specifically, we define protoclusters as the ensemble of galaxies that will reside within $R_{\rm{200c}}$ at $z=0$. To ensure sufficient stellar mass resolution, we include only galaxies, regardless of their redshift, that have reached a minimum stellar mass of ${M}_{\star}\geq 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. This choice implies that the number of protocluster members will generally increase as redshift decreases, with higher-redshift protoclusters having fewer members that meet the minimum stellar mass threshold.
Observationally, a wide range of stellar mass completeness limits exist due to the heterogeneous sampling of protoclusters. These samples are typically detected using selection functions that favor massive, star-forming galaxies, and are thus generally limited to galaxies at the massive end of the stellar mass function (${M}_{\star}>10^{10}\leavevmode\nobreak\ {\rm M}_{\odot}$). However, protoclusters identified in photometric and spectroscopic studies have achieved stellar mass completeness down to ${M}_{\star}\gtrsim 10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ (Lemaux et al., 2022). Consequently, the protocluster galaxies explored in this study are at least an order of magnitude less massive than those in the most complete observational samples to date.
Fig. 2 shows the spatial distribution of protocluster galaxies at different redshifts, organized by $z=0$ cluster mass: massive clusters (first and second rows), intermediate-mass clusters (third and fourth rows), and low-mass clusters (fifth and sixth rows). Black circles indicate galaxies with ${M}_{\star}\geq 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ (red circle) at $z=0$, forming our “baseline” protocluster population. Gray circles represent galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside between $R_{\rm{200c}}<R<10\leavevmode\nobreak\ R_{\rm{200c}}$ at $z=0$. From left to right, the columns show the spatial distribution of these galaxies across six snapshots from $z=0$ to $z=5$. A defining characteristic of protoclusters is their vast spatial extent, spanning several physical megaparsecs (or tens of comoving megaparsecs) at early cosmic times. This figure highlights a major observational challenge: obtaining high-quality spectroscopic redshifts over a sufficiently large area to accurately distinguish galaxies that will eventually belong to a virialized cluster at $z=0$ from those that will not. To further explore this, in Appendix A we examine the observational feasibility of the protocluster definition used in this work, specifically quantifying contamination from galaxies that may be identified as part of a protocluster but will not reside within the cluster’s virial radius at $z=0$.
2.4 Characterizing Protocluster Sizes
As illustrated in Fig. 2, galaxy protoclusters are extremely extended, with sizes exceeding tens of comoving megaparsecs. To characterize the inner and outer regions of the simulated protoclusters explored in this study we adopt two characteristic radii, $R_{10}$ and $R_{90}$, which, at a given redshift, represent the radii enclosing $10\%$ and $90\%$, respectively, of the total stellar mass of galaxies with ${M}_{\star}\geq 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ at $z=0$.
This definition is most similar to that employed by Muldrew et al. (2015), who defined $R_{90}$ as the radius enclosing $90\%$ of the stellar mass of their simulated protoclusters, which were sourced from the Millennium Simulation (Springel et al., 2005) with the Guo et al. (2011) semi-analytic model applied. However, unlike in Muldrew et al. (2015), where $R_{90}$ was measured with respect to the center of the galaxy cluster at $z=0$, we measure $R_{10}$ and $R_{90}$ with respect to the center of mass of the protocluster at a given redshift.
Our definition provides a metric that is more accessible to observers, as the center of mass of observed galaxy protoclusters can be estimated. Nevertheless, on average, our inferred $R_{90}$ and its redshift evolution are consistent with those of Muldrew et al. (2015), in the scenario where protocluster member galaxies are defined as progenitors of galaxies residing within $R_{\rm{200c}}$ at $z=0$.
2.5 Constructing Protocluster Galaxy Density Maps
Observationally, the local environment of galaxy clusters and protoclusters has been characterized using a variety of methods that estimate galaxy density in three-dimensional space (RA, Dec, and redshift), such as nearest neighbors (e.g., Polletta et al., 2021; Champagne et al., 2021), Friends-of-Friends algorithms (e.g., Calvi et al., 2021; Helton et al., 2024a), tessellation-based density estimators (e.g., Ramella et al., 2001; Cooper et al., 2005; Cucciati et al., 2014; Darvish et al., 2015; Lemaux et al., 2018; Hung et al., 2020; Sarron & Conselice, 2021; Forrest et al., 2023), and Gaussian kernel density estimation (KDE) (e.g., Bădescu et al., 2017; McConachie et al., 2022; Brinch et al., 2023; McConachie et al., 2025). To probe the local environment of the simulated protoclusters in this study, we use adaptive binning with nearest-neighbor weighting. Specifically, we construct a three-dimensional grid using the $x$, $y$, and $z$ positions of individual galaxies from a given redshift snapshot, with a uniform bin size in comoving Mpc (cMpc) that spans the entire protocluster. The grid is constructed in cMpc, but for visualization purposes the density maps are scaled to physical Mpc (pMpc) to better capture the protocluster environment at different redshifts, given the smaller dynamic range in physical units. Moreover, while the bin size is fixed at given redshift, we adaptively adjust it to account for protocluster size evolution, setting it to $0.5$ cMpc for $10<z<$ 1, $0.75$ cMpc for $41\leq z<$ 4, and $1$ cMpc for $z\geq 4$.
Using this grid, we create a multi-dimensional histogram to count galaxies in each cell and measure galaxy number density. To incorporate the local galaxy environment, we calculate the three-dimensional distance to the nearest neighbor for each galaxy and apply an exponential decay weighting, where the weight decreases with increasing neighbor distance. Specifically, the weight for each galaxy is $w=\exp(-d_{\mathrm{nn}}/\lambda_{\mathrm{nn}})$, where $\lambda_{\mathrm{nn}}$ is a parameter controlling the rate of decay, which we fix to a value of $10$ cMpc. This approach yields results quantitatively similar to the KDE approach but with significantly lower computational cost.
The final density map is generated by applying a Gaussian filter to the weighted 3D histogram and summing galaxy densities along the $z$-axis to create a 2D projected density field. The highest-density region is identified as the maximum density value within this field. To locate it, we first find the maximum density in the $x$-$y$ plane and then trace along the $z$-axis to the cell with the highest galaxy concentration. Finally, the highest-density peak is defined as the centroid of the galaxies in this region.
Figure 3: Top: Protocluster size evolution in physical units, characterized by $R_{90}$ and $R_{10}$, which represent the radii, measured relative to the center of mass of the protocluster at the observed redshift, that enclose $90\%$ and $10\%$, respectively, of the total stellar mass of the baseline protocluster population, defined as galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ at $z=0$. The purple, blue, and green lines indicate the mean values in bins of $z=0$ halo mass, with bands showing the corresponding $68\%$ quantiles. Middle: Protocluster size evolution in comoving units. Bottom: Protocluster size evolution in units of arcminutes. Open markers show the angular sizes of spectroscopically-confirmed protoclusters from the literature with $>10$ members at $52<z<$ 5. Sizes correspond to the aperture radius or half the field of view; for asymmetric fields, we use half the average of both dimensions. Data and references are in provided in Table 1 of Appendix B. These results suggest that the vast majority of existing spectroscopically-confirmed protoclusters do not fully capture the expected volume occupied by the progenitors of the most massive galaxy clusters.
Figure 4: The projected galaxy density distribution at $z=3$ for 15 protoclusters that will collapse into clusters by $z=0$. The density maps are constructed using adaptive binning with nearest-neighbor weighting, smoothed with a Gaussian filter, summed along the $z$-axis, and visualized using filled contours. The rows depict the projected galaxy density distribution for the progenitors of the five most massive clusters at $z=0$ (top row), five intermediate-mass clusters (middle row), and the five least massive clusters at $z=0$ (bottom row). Individual galaxies are shown as white dots, with their sizes scaling with stellar mass. Here the galaxies represent the “baseline” population, defined as galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ at $z=0$. We highlight four regions of interest: the center of mass (red X), the highest galaxy density peak (red cross), the most massive galaxy at the given redshift (red square), and the progenitor of the $z=0$ BCG (red star). We also overplot $R_{10}$ for each protocluster, illustrating how this radial extent varies among protoclusters at this epoch. A key takeaway is that the location of the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$) does not always correspond to the location of the most massive galaxy or the BCG progenitor.
Figure 5: The projected galaxy density distribution as a function of redshift for the progenitor of the most massive galaxy cluster at $z=0$. The construction of these density maps follows the same procedure outlined in Fig. 4 and §2.5. The top row shows the galaxy density maps measured using the baseline protocluster population, defined as galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ at $z=0$. This serves as a point of comparison for the middle and bottom rows, which show the galaxy density maps for observationally limited protocluster galaxy samples. Namely, the middle row is limited to a subsample of the baseline population with ${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, whereas the bottom row is a subsample with $\mathrm{SFR}>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. Key regions are highlighted at each snapshot: the center of mass (red), the baseline highest galaxy density peak (red cross), the observationally limited highest galaxy density peak (red plus), the most massive galaxy at the given redshift (red square), and the progenitor of the $z=0$ BCG (red star). We overplot $R_{10}$ for each protocluster as a dashed red circle, illustrating how this radial extent evolves with redshift. For this particular cluster, we find the ${M}_{\star}$-limited sample generally fails to recover the baseline highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$) at $z>2$, while the SFR-limited subpopulation struggles at $z>4$.
Figure 6: Three-dimensional separation between the baseline highest galaxy density peak and four regions of interest as a function of redshift: (i) the most massive galaxy at a given redshift (top-left), (ii) the BCG progenitor at a given redshift (top-right), (iii) the ${M}_{\star}$-limited (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) highest density peak (bottom-left), and (iv) the SFR-limited (SFR $>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) highest density peak (bottom-right). The separations are binned by cluster halo mass at $z=0$, with the purple, blue, and green lines representing the median separations for progenitors of massive, intermediate, and low-mass clusters, respectively, and the error bars showing the 16th to 84th percentiles. While there is significant protocluster-to-protocluster variation in the highest density peak separation, in general, the baseline highest galaxy density peak does not coincide with the most massive galaxy or BCG progenitor at $z>1$. Likewise, at $z>2$, the highest galaxy density peaks identified by ${M}_{\star}$- and SFR-limited subpopulations generally do not trace the baseline highest galaxy density peak.
Figure 7: Cumulative recovery fraction of the baseline highest galaxy density peak as a function of the three-dimensional separation between the baseline highest galaxy density peak and the four regions of interest highlighted in Fig. 6. Solid lines show the cumulative recovery fractions in redshift bins from $z=1$ to $z=5$. In all cases at $z>2$, the recovery fractions are low, typically $\lesssim 35-40\%$ within an accuracy of 1.0 pMpc. This indicates that the most massive galaxy and BCG progenitors, as well as the highest galaxy density peaks traced by the ${M}_{\star}$- and SFR-limited subpopulations, do not typically coincide with the baseline highest galaxy density peak on scales smaller than 1 pMpc at $z>2$.
3 Results
3.1 Comparing Sizes of Observed and Simulated Protoclusters
In the top and middle panels of Fig. 3 we present the redshift evolution of $R_{10}$ and $R_{90}$ in physical and comoving megaparsecs. To explore how these sizes depend on cluster mass at $z=0$, we divide the data into three mass bins: low-mass ($10^{14.2}<M_{\rm{200}}^{z=0}/{\rm M}_{\odot}<10^{14.6}$), intermediate-mass ($10^{14.6}<M_{\rm{200}}^{z=0}/{\rm M}_{\odot}<10^{15.0}$), and high-mass ($10^{15.0}<M_{\rm{200}}^{z=0}/{\rm M}_{\odot}<10^{15.5}$) clusters. The solid and dashed lines represent the average values of $R_{10}$ and $R_{90}$, respectively, while the shaded bands indicate the 16th to 86th percentile range. Consistent with Muldrew et al. (2015), which adopts a similar definition of protocluster size, we find that progenitors of more massive clusters at $z=0$ are substantially more extended at earlier times, while those of lower-mass clusters are more compact, having sizes approximately $\sim 40\%$ smaller than their massive counterparts at $z\gtrsim 3$.
[…]
In the bottom panel of Fig. 3 open markers indicate the angular sizes of spectroscopically-confirmed protoclusters from the literature, with redshifts in the range $52<z<$ 5 and each containing at least $10$ spectroscopic members. These sizes correspond to either the aperture radius used to define the protocluster or half the field of view. When the field of view has unequal length and width, we use half of the average of the two dimensions. The only exception is the study by Shah et al. (2024), where the sizes are derived from the protocluster volumes provided in their work. This data, along with the corresponding references, are provided in Table 1 of Appendix B.
This comparison highlights that most spectroscopically-confirmed protoclusters in the literature do not probe the full expected volume of the progenitors of the most massive clusters at $z=0$. Instead, most of these studies are limited to aperture sizes that correspond to the protocluster core, or they lie in a region of confusion around $\sim 4\arcmin$, where they could be probing either the core of a massive cluster progenitor or the full extent of a low-mass cluster progenitor. As a result the inferences drawn from these spectroscopically-confirmed protocluster populations may be misleading or incomplete, as these surveys lack the necessary field of view and sensitivity to capture the bulk of the underlying protocluster population at these redshifts, especially for the most massive cluster progenitors.
3.2 3D Separation of Galaxy Density Peaks: Complete vs. Observationally-limited Populations
Fig. 4 shows the projected galaxy density distribution at $z=3$, measured relative to the center of mass, for the “baseline population”, defined in §2.3 as all galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that by $z=0$ will reside within $R_{\rm{200c}}$. As detailed in §2.5 this galaxy density map is generated using adaptive binning with nearest-neighbor weighting. The top, middle, and bottom rows display the projected galaxy density distribution for the protoclusters that will collapse into the five most massive (top row), five least massive (bottom row), and five intermediate-mass (middle row) clusters. In each panel we highlight key locations such as the highest galaxy density peak (red cross), the most massive galaxy (red square), the center of stellar mass (red X), and the $z=0$ brightest cluster galaxy (BCG) progenitor (red star). We also overplot $R_{10}$ centered on the highest galaxy density peak, but measured at the location of the center of mass, as described in §3.1. Finally, the white circles show the projected locations of the galaxies that make up the protocluster, with their sizes scaled according to their stellar mass.
A key takeaway from Fig. 4 is that at $z=3$ the location of the highest galaxy density peak does not always correspond to the location of the most massive galaxy or the BCG progenitor. This highlights the diverse evolutionary states of protoclusters at a given redshift. Moreover, it suggests that the region of a protocluster with the highest galaxy concentration is not necessarily traced by the most massive galaxies. While the highest density peaks shown here are measured for protocluster populations with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, current observational surveys do not achieve this level of stellar mass completeness over the volumes spanned by protoclusters at $z=3$. Since current spectroscopically-confirmed protoclusters under-sample the faint end of the galaxy luminosity function, the inferred galaxy density distributions – and consequently, the highest galaxy density peaks – may not accurately reflect the underlying distribution. In other words, the true highest galaxy density region may be missed entirely by observed protocluster samples.
In Fig. 5 we examine whether the highest galaxy density peak, as measured using our baseline protocluster population, would be misidentified if the protocluster population were limited to massive or highly star-forming galaxies. Specifically, we plot the projected galaxy density distribution at five redshift snapshots from $z=1$ to $z=5$ for the progenitor of the most massive cluster at $z=0$. The top row shows the projected galaxy density distribution for the baseline population, while the middle and bottom rows show the stellar mass- and SFR-limited subpopulations, defined by galaxies with a minimum stellar mass of $>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ or $\mathrm{SFR}>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$, respectively.
For this cluster, we observe a discrepancy between the locations of the highest galaxy density peaks traced by the baseline population and the stellar mass-limited subpopulation at $z>2$, whereas for the SFR-limited subpopulation, the discrepancy occurs at $z=1$ and $z=5$. At $z=5$ the stellar mass- and SFR-limited subpopulations fail to recover the baseline density peak primarily because their selection functions exclude fainter and lower-mass protocluster members. However, at $z=1$, the SFR-limited subpopulation fails to recover the baseline density peak as the underlying cluster population becomes increasingly dominated by quiescent galaxies.
While visualizing the separation between the baseline and stellar mass/SFR-selected highest density peaks is valuable, quantifying these separations as a function of redshift is more insightful. This is shown in Fig. 6, which shows the three-dimensional separation between the location of the highest galaxy density peak for the baseline protocluster population and other key locations: the most massive galaxy (top-left panel), the BCG progenitor (top-right panel), the SFR-limited ($>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) highest galaxy density peak (bottom-right panel), and the stellar mass-limited ($>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) highest galaxy density peak (bottom-left panel). The purple, blue, and green lines represent the median results binned by cluster halo mass at $z=0$, with error bars indicating the 16th and 84th percentile range.
In comparing the separation between the location of the baseline highest galaxy density peak and the location of the most massive galaxy, we find that above $z>1$ these two locations show very little overlap. The magnitude of this separation correlates modestly with the cluster halo mass at $z=0$. Similarly, the separation between the baseline highest galaxy density peak and the BCG progenitor shows little agreement beyond $z>1$, with the degree of separation again correlating with the mass of the cluster at $z=0$. In both cases, there is significant protocluster-to-protocluster variation, indicated by the 1-$\sigma$ spread in highest-density peak separations at fixed redshift, largely driven by differences in protocluster evolutionary states.
For the stellar mass-limited subpopulation (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$), the separation between the highest galaxy density peak and the baseline peak exceeds $1$ pMpc at $z>2$, showing a relatively milder dependence on the cluster halo mass at $z=0$. The most striking difference occurs when comparing the baseline highest galaxy density peak to the peak probed by the SFR-limited subpopulation ($\mathrm{SFR}>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$). Despite sharing the same stellar mass limit (${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$), we find that the overlap is poor at all redshifts. This suggests that observations focusing on highly star-forming galaxies will miss the highest concentration of galaxies during the protocluster stage of cluster assembly.
In Fig. 7, we show the cumulative recovery fraction of the baseline highest galaxy density peak as a function of the three-dimensional separation from this peak to four key locations: the most massive galaxy (top-left panel), the BCG progenitor (top-right panel), and the highest galaxy density peaks traced by the ${M}_{\star}$- and SFR-limited populations (bottom-left and right panels, respectively). The recovery fraction is defined as the ratio of samples that recover the baseline highest galaxy density peak within a given separation to the total number of samples. As indicated by the five colored lines in Fig. 7, the results are presented as cumulative sums across five redshift snapshots from $z=1$ to $z=5$. In general, the baseline highest galaxy density peak is only recovered in $\lesssim 35-40\%$ of cases within an accuracy of 1.0 pMpc ($\sim 2-2.6\arcmin$) at $z>2$. In other words, the most massive galaxies, BCG progenitors, and the highest galaxy density peaks traced by the ${M}_{\star}$- and SFR-limited populations (constrained to ${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ and SFR$>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$, respectively) do not typically coincide with the true highest galaxy density peak on scales smaller than 1 pMpc at $z>2$.
Figure 8: Three-dimensional separation as a function of redshift between the baseline highest galaxy density peak and peaks inferred from ${M}_{\star}$- and SFR-limited subsamples (left and right columns, respectively). Rows correspond to protoclusters binned by halo mass at $z=0$, with progenitors of the most (least) massive clusters shown in the top (bottom) row. The highest galaxy density peaks in the ${M}_{\star}$-limited subsamples are identified using stellar mass thresholds ranging from $10^{9.0}$ to $10^{10.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. Similarly, the SFR-limited peaks are selected using SFR thresholds from $0.1$ to $5.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$, with an additional stellar mass restriction of ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {\rm M}_{\odot}$. While there is substantial protocluster-to-protocluster variation in peak-centric separations, as indicated by the 1-$\sigma$ error bars, both the ${M}_{\star}$- and SFR-limited subpopulations generally struggle to recover the baseline highest density peak at $z>2$, with median separations exceeding 1 pMpc even for the least restrictive thresholds.
[…]
3.3 Evaluating Completeness Limits for Recovering the True Highest Galaxy Density Peak
In Fig. 8 we explore the stellar mass and SFR completeness limits required to recover the location of the baseline highest galaxy density peak. The left column presents the 3D separation between the baseline and stellar mass-limited highest galaxy density peaks, with the top, middle, and bottom rows showing the data binned by cluster halo mass at $z=0$. Separations are plotted as a function of redshift, with three stellar mass thresholds, ranging from $10^{9.0}$ to $10^{10.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. While the scatter is significant, the separations generally correlate with the cluster halo mass at $z=0$, with the median separation increasing for more massive $M_{\rm{200}}^{z=0}$ bins. We also find that separations are sensitive to changes in the minimum stellar mass, as even a $0.5$ dex increase from the baseline stellar mass (i.e., ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) results in separations $\gtrsim 1$ pMpc at $z\geq 3$.
In the right column of Fig. 8, we show the 3D separation between the baseline and SFR-limited highest galaxy density peaks, with the SFR-limited sample restricted to ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {\rm M}_{\odot}$ and ranging from $0.1$ to $5.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. In general, we find that the correlation between the cluster halo mass at $z=0$ and the magnitude of the median separations is relatively weak. Additionally, while the scatter is large, the median separations generally exceed $0.5$ pMpc at $z>2$ and increase with more restrictive SFR thresholds.
In Fig. 9, we plot the baseline highest density peak recovery fraction as a function of highest density peak-centric separation. The left column shows the cumulative recovery fraction, binned by redshift from $z=1$ to $z=5$, for the separation between the baseline highest galaxy density peak and ${M}_{\star}$-limited highest galaxy density peaks for stellar mass thresholds ranging from $10^{9.0}$ to $10^{10.5}\leavevmode\nobreak\ {\rm M}_{\odot}$. For ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {M}_{\star}$, approximately $40\%$ of the sample recovers the baseline density peak for separations less than $1$ pMpc; however, this drops to $20\%$ for peaks traced by protocluster populations with ${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$.
The right column of Fig. 9 shows the cumulative recovery fraction for the separation between the baseline highest galaxy density peak and the highest galaxy density peaks traced by SFR-limited subpopulations with SFR thresholds ranging from $0.1$ to $5.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. For all of the SFR thresholds explored, the recovery rate is consistently around $40\%$ at $z>2$ for scales less than 1 pMpc. Given that this sample is constrained to ${M}_{\star}>10^{9.0}\leavevmode\nobreak\ {\rm M}_{\odot}$, it suggests that stellar mass, rather than SFR, is the primary driver of the discrepancy between the locations of the observationally-limited and baseline highest galaxy density peaks. Overall, these results show that the location of the highest concentration of galaxies for protocluster populations complete down to ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ is generally not recovered on scales less than 1 pMpc at $z>2$.
Figure 10: Ratio of the average SFR (top row) and stellar mass (bottom row) within $R_{10}$ centered on the baseline highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$), compared to the average SFR and stellar mass in the surrounding volume outside of $R_{10}$. The columns show these ratios binned by cluster mass at $z=0$. A box-and-whisker plot visualizes the distribution, with vertical boxes representing the interquartile range (25th to 75th percentiles), horizontal bars and circles indicating the median and mean ratios, and whiskers extending to 1.5 times the interquartile range, covering approximately $95\%$ of the distribution. While the mean ratios are generally above unity, they are sensitive to outliers. The median ratios, which are robust to outliers, show that the average SFR and stellar mass near the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$) are lower than in the protocluster as a whole. This suggests that regions with the highest galaxy densities are not necessarily associated with enhanced star formation or stellar mass growth.
Figure 11: Projected density distribution for the progenitor of the most massive $z=0$ galaxy cluster as a function of redshift. The top row shows the galaxy density map, while the middle and bottom rows display the stellar mass and dark matter weighted density maps from TNG-Cluster, respectively. The location of the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$) is marked by a black cross, while the highest stellar mass ($\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$) and dark matter ($\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$) peaks are indicated by black plus and hexagon symbols, respectively. In this cluster the locations of these highest density peaks overlap at $z\leq 2$ but are separated by $1-3$ pMpc at earlier times. This suggests that the highest galaxy density peak does not necessarily coincide with the highest concentrations of stellar mass and dark matter in protoclusters.
3.4 Comparing Galaxy Properties Inside vs. Outside Highest Galaxy Density Peaks
To examine the impact of potentially missing the highest galaxy density peak in observations, we explore the properties of galaxies (e.g., SFRs or stellar masses) inside and outside this region in our simulated protocluster population. We compare these properties by calculating the ratio of the average SFR and stellar mass inside and outside $R_{10}$, measured relative to the location of the highest galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$). Ratios close to unity imply that the galaxies in the highest density peak have properties on average similar to those of galaxies in the rest of the protocluster. Ratios below or above unity indicate that the galaxies within the highest density peak have properties that are either depressed or elevated relative to the remainder of the protocluster.
In Fig. 10 we show the ratio of the average SFR and stellar mass as a function of redshift for galaxies within $R_{10}$ centered on the highest galaxy density peak, compared to those beyond this region. The data is binned by cluster halo mass at $z=0$ and presented as a box-and-whisker plot, with the vertical boxes representing the interquartile range (25th to 75th percentiles), and the horizontal bars and circles showing the median and mean of the distribution. The whiskers extend to 1.5 times the interquartile range, encompassing approximately $95\%$ of the distribution.
The top row of Fig. 10 shows that for the average SFR within the highest galaxy density peak the mean of this distribution is consistently above unity, regardless of the cluster mass at $z=0$, compared to the remainder of the protocluster. As this trend can be influenced by outliers, we also examine the median of the distribution. We find that, regardless of the cluster mass at $z=0$, the median SFR within the highest galaxy density peak is below unity for $z>2$, indicating that the average SFR in this region is depressed relative to the rest of the protocluster.
In the bottom row of Fig. 10 we show the ratio of the average stellar mass within the highest density peak compared to the average stellar mass of the protocluster. The mean of this distribution is consistently above unity for all redshifts, with very little dependence on the halo mass of the cluster at $z=0$. Meanwhile, we find that the median stellar mass ratio is generally below unity at $z>3$ but increases with redshift toward $z=0$. Additionally, we observe a mild dependence on the host halo mass of the cluster at $z=0$, with progenitors of the most massive clusters reaching unity earlier. These results imply that the regions with the highest concentration of galaxies in protocluster populations complete down to ${M}_{\star}\sim 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ are typically not regions of accelerated galaxy evolution relative to the remainder of the protocluster.
Figure 12: Cumulative recovery fraction of the baseline stellar mass and dark matter highest density peaks, binned by redshift and measured relative to the highest galaxy density peaks traced by various ${M}_{\star}$- and SFR-limited subpopulations. The first two rows show separations from the baseline stellar mass (first row) and dark matter (second row) highest density peaks to the highest galaxy density peak traced by ${M}_{\star}$-limited subpopulations. The third and fourth rows show the same for SFR-limited subpopulations. Columns represent increasingly restrictive stellar mass ($10^{8.5}$ to $10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$) and SFR (0.10.10.10.$10.0\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$) thresholds. Regardless of the selected ${M}_{\star}$ or SFR threshold, the highest stellar mass and dark matter density peaks are recovered in $\lesssim 40\%$ of cases with accuracies within 1 pMpc ($2.1-2.6\arcmin$) for $2<z\leq 5$. However, in the vast majority of cases, these peaks are only recovered at larger separations of 14141–41 – 4 pMpc.
3.5 Measuring 3D Separation of Matter Density Peaks: Complete vs. Observationally-limited Populations
The results presented in the previous section appear to conflict with findings from observed protocluster studies, which suggest that the densest regions of galaxy protoclusters are sites of accelerated galaxy growth and prodigious star formation (e.g., Steidel et al., 2005; Dannerbauer et al., 2014; Hayashi et al., 2016; Miller et al., 2018; Shimakawa et al., 2018a). However, this discrepancy may stem from the specific region of the galaxy protocluster being probed. Thus far we have considered only the region of the highest galaxy density peak, which is commonly used to characterize the density field of observed protocluster galaxies. Alternatively, the highest matter density peak could be examined by weighting the galaxy density field according to the stellar and halo masses of the protocluster galaxies. To achieve a balanced weighting scheme that prevents either high-mass or low-mass galaxies from dominating the density field, we define the weights applied to the 3D histogram as $w=M^{\alpha}+C$, where $M$ represents either the stellar mass or dark matter component of each galaxy. We set $\alpha=0.5$ to ensure a balanced contribution from low-mass and high-mass galaxies, while the constant $C=0.5$ establishes a baseline weight [^2] In practice, $\alpha$ only affects the map in the range between 0 and 1, with $\alpha\leq 0$ returning the original galaxy density map, and $\alpha\geq 1$ excluding contributions from low-mass galaxies, making the map solely reflect the location of the most massive galaxy..
In Fig. 11 we present the projected galaxy density (top row), stellar mass density (middle row), and halo mass density (bottom row) for the progenitor of the most massive galaxy cluster at $z=0$. The columns display the projected density maps at five distinct redshifts, ranging from $z=1$ to $z=5$. We overlay the location of the highest density peak for each density field, marking the highest galaxy density peak with a black cross, the highest stellar mass peak with a black plus sign, and the highest dark matter mass density peak with a black hexagon. For this specific test case we find that above $z>2$ the location of the highest galaxy density peak does not always coincide with the location of the highest stellar mass or halo mass density peaks. Observationally, the opposite is typically observed, for instance, Koyama et al. (2021) and Polletta et al. (2021) find that the densest region of the Planck-selected protocluster PHz G237.01+42.50 contains the most massive members.
Figure 13: Box-and-whisker plot of the ratio of the average SFR (top row) and stellar mass (bottom row) within $R_{10}$ centered on the highest density peak for dark matter (blue boxes, $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$), stellar mass (orange boxes, $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$), and galaxies (yellow boxes, $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$), compared to the average SFR and stellar mass in the surrounding volume outside $R_{10}$. The columns display these ratios in bins of cluster mass at $z=0$. In contrast to the ratios for $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$, the matter density peaks exhibit ratios significantly above unity for both SFR and stellar mass, suggesting that galaxies in these regions experience accelerated growth and enhanced star formation compared to the protocluster as a whole.
In Fig. 12, we plot the cumulative recovery fraction of the baseline highest matter density peaks as a function of their relative separation from the highest galaxy density peaks, binned by redshift. These peaks are traced by a total of eight ${M}_{\star}$- and SFR-limited subpopulations, including galaxies from all 352 protoclusters. The recovery fraction is defined as the ratio of samples that recover the baseline density peak within a given separation to the total number of samples.
The first two rows of Fig. 12 show the cumulative matter density recovery fractions measured from the baseline stellar mass (first row) and dark matter (second row) highest density peaks to the highest galaxy density peak traced by four ${M}_{\star}$-limited subpopulations. Moving from left to right, the stellar mass thresholds increase from $10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ (i.e., the baseline stellar mass established in this study) to $10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$. Regardless of the stellar mass threshold, the true highest stellar mass and dark matter density peaks are only recovered in $\lesssim 40\%$ of cases within an accuracy of 1.0 pMpc ($\sim 2.1-2.6\arcmin$) at $2<z\leq 5$. However, at $z=5$, the recovery fraction of the highest stellar mass density peak increases with more restrictive stellar mass thresholds when considering separations within 0.5 pMpc ($1.3\arcmin$). Specifically, the recovery fraction rises from $15\%$ for ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ to $30\%$ for ${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$, suggesting that samples containing more massive galaxies better trace the highest stellar mass density peak on small scales.
A similar trend, but in the opposite direction, is observed for the recovery fraction of dark matter density peak at $z=5$ when considering separations with accuracies within 0.5 pMpc ($1.3\arcmin$). In this case, the recovery fraction decreases from $30\%$ for ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ to $25\%$ for ${M}_{\star}>10^{10.0}\leavevmode\nobreak\ {\rm M}_{\odot}$. Nevertheless, in most cases, the matter density peaks are only recovered at separations between 1 and 4 pMpc, suggesting that galaxy density peaks inferred from ${M}_{\star}$-limited populations do not generally trace the true matter density peaks.
The last two rows of Fig. 12 show the cumulative matter density recovery fractions measured from the baseline stellar mass (third row) and dark matter (fourth row) highest density peaks to the highest galaxy density peak traced by four SFR-limited subpopulations. Moving from left to right, the SFR thresholds increase from $0.1$ to $10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr}^{-1}$. The SFR-limited results, which are not additionally constrained by stellar mass as in Fig. 9, are quite similar to the stellar mass-limited results. Specifically, matter density peaks are recovered in only $\sim 40\%$ of cases within an accuracy of 1.0 pMpc ($\sim 2.1-2.6\arcmin$) for $2<z\leq 5$, while the majority of samples recover the highest matter density peaks only at separations between 1 and 4 pMpc for $z>2$.
The cumulative recovery fraction results from Fig.12 suggest that regardless of the stellar mass or SFR limitations of the protocluster galaxy population, the true matter density peaks are only recovered in $\sim 40\%$ of cases with an accuracy of 1.0 pMpc ($\sim 2.1-2.6\arcmin$) for $2<z\leq 5$. However, these results depend on the weighting scheme used in the matter density maps. The chosen value of $\alpha=0.5$ provides a balanced approach, with $\alpha\leq 0$ reproducing the baseline galaxy density field and $\alpha\geq 1$ being overly influenced by the location of the most massive galaxy. Slight deviations from this parameter choice would only marginally affect the agreement between the galaxy and matter density peaks. Despite this caveat, these results suggest that the highest density peaks traced by galaxies do not always coincide with the location of matter density peaks, likely due to the fact that the baseline highest galaxy density peak does not always coincide with the most massive galaxies or BCG progenitors, which are generally more reliable tracers of the matter density field (see Fig. 6).
3.6 Comparing Galaxy Properties Inside vs. Outside Highest Matter Density Peaks
To compare galaxies in the highest matter density peaks to those outside these regions, we compute the ratio of average stellar mass and SFR inside and outside $R_{10}$, centered on these peaks. Fig. 13 shows the ratio of average SFR (top) and stellar mass (bottom) versus redshift for galaxies within $R_{10}$, centered on the highest dark matter density peak ($\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$), stellar mass density peak ($\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$), and galaxy density peak ($\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$). Although not shown, we also examined the total baryonic mass density (stars + gas) and found it closely follows the dark matter density field. From left to right, the columns bin the ratios by protoclusters that will evolve into high-, intermediate-, and low-mass clusters by $z=0$. Fig. 13 presents the data as a box and whisker plot, with blue, orange, and yellow boxes showing the ratio distributions relative to $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$$\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$, and $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$, respectively. Horizontal bars and circles indicate the median and mean of each distribution.
The top row of Fig. 13 shows that, whether probed by the mean or the median, the average SFR centered on $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$ and $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$ is significantly above unity for $z>1$. This contrasts sharply with the results relative to $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$, where the median of the distribution is below unity for $z>2$. At $z=1$, all three distributions return comparable ratios of the average SFR within and beyond the respective highest density peak, reflecting the results from the top-left panel of Fig. 12, where the locations of these density peaks overlap by this epoch.
Similarly, the bottom row of Fig. 13 shows that, whether assessed by the mean or the median, the typical stellar mass within and beyond $\vec{x}_{\delta_{\mathrm{DM,peak}}}^{\mathrm{Baseline}}$ and $\vec{x}_{\delta_{\mathrm{\star,peak}}}^{\mathrm{Baseline}}$ is also significantly above unity at $z\geq 2$, with earlier epochs showing the strongest enhancement. Once again, this contrasts sharply with the median of the distribution of the ratio of the average stellar mass within and beyond $\vec{x}_{\delta_{\mathrm{gal,peak}}}^{\mathrm{Baseline}}$, which is below unity for $z>2$. However, by $z=1$ all three distributions return comparable ratios of the average stellar mass within and beyond the respective highest density peak.
The results in Fig. 13 show that without weighting by stellar mass, the highest density peaks traced by galaxy populations complete down to ${M}_{\star}=10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ do not generally correspond to regions of accelerated mass growth or enhanced star formation relative to the rest of the protocluster. This suggests that observed protocluster density maps, which are more heavily weighted towards more massive galaxies due to the incompleteness of lower-mass galaxies (e.g., ${M}_{\star}<10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) at $z>2$, may overestimate how often the inferred highest density regions are sites of accelerated galaxy evolution.
4 Discussion
4.1 Implications of Misidentifying the True Highest Galaxy Density Peak in Protoclusters
[…]
The misidentification of the highest galaxy density peak has important consequences. When comparing galaxies near the highest galaxy density peak to those in the broader protocluster, we find that galaxies in these peaks have, on average, lower SFRs and stellar masses relative to those outside, seemingly contradicting previous studies that suggest the densest regions of protoclusters are sites of accelerated mass growth and enhanced star formation. This discrepancy arises because the highest galaxy density peaks do not necessarily coincide with the highest matter density peaks. Instead, the latter — traced by stellar mass and dark matter — exhibit the expected signs of rapid mass growth and high SFRs. Similarly, Muldrew et al. (2018) found that the largest progenitor halos in protoclusters host the highest stellar mass and SFR densities, attributing this to these halos being biased tracers of the underlying dark matter density field (Tinker et al., 2010). This suggests that observational studies, often limited to massive, star-forming protocluster galaxies, are prone to overestimating how frequently the inferred highest-density regions are sites of accelerated stellar mass growth and enhanced star formation, as they likely miss the true highest-density regions.
If future observations probe lower-mass galaxies (e.g., ${M}_{\star}\gtrsim 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$), they may find that the highest galaxy density regions are not necessarily sites of enhanced galaxy evolution compared to the broader protocluster population. However, it is unclear whether these galaxies will exhibit enhanced SFRs and accelerated mass growth compared to a coeval field population. This would be an interesting avenue for exploration, as current observational evidence suggests that while the densest regions in protoclusters show star formation enhancement relative to the field, the majority of protocluster members are on the main sequence (Shimakawa et al., 2018b; Polletta et al., 2021; Pérez-Martínez et al., 2023, 2024a, 2024b). However, Hayashi et al. (2016) found that relatively low-mass galaxies (${M}_{\star}\gtrsim 10^{9.3}\leavevmode\nobreak\ {\rm M}_{\odot}$) in the USS 1558-003 protocluster at $z=2.5$ exhibit higher SFRs than main-sequence galaxies at this redshift. These results highlight the need to expand the stellar mass and SFR limits of observational samples to determine if these trends extend to unseen members and to acquire large, spectroscopically-confirmed protocluster samples to assess how variance in protocluster evolutionary states influences these trends.
4.2 Considerations for Observational Studies of Galaxy Protoclusters
Our analysis suggests a few key considerations for observational studies of galaxy protoclusters. First, aperture sizes smaller than $4\arcmin$ should be avoided, as they fail to capture the full extent of the progenitors of the most massive galaxy clusters. These apertures also probe a region of confusion, as they can overlap with the innermost regions of the progenitors of the most massive clusters and the full extent of the progenitors of the lowest mass clusters (see Fig. 2). Instead, apertures larger than $8\arcmin$ are recommended, as they will capture the full extent of the progenitors of the most massive clusters at $z>2$.
[…]
Finally, the incompleteness in current protocluster selection functions, both in terms of stellar mass and SFR, inherently biases observed density maps. As shown in Fig. 13, the identification of the highest galaxy density peaks as regions of enhanced star formation and stellar mass growth depends on how the maps are weighted. While it is difficult to fully correct for this missing data, one approach would be to analyze analogs of the observed protoclusters at similar redshifts in cosmological simulations. Information provided by these models, such as the abundance, spatial distribution, and clustering of galaxies below the SFR and stellar mass sensitivity limits of the observations, could inform adjustments to density maps.
5 Summary and Conclusions
In this paper we leverage the unique combination of a large sample of very massive galaxy clusters (and their progenitors) and high baryonic mass resolution of the TNG-Cluster simulation to examine how observational incompleteness impacts inferences about galaxy protocluster populations. We achieve this by first defining our protocluster galaxy population as the ensemble of galaxies with ${M}_{\star}>10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$ that will reside within $R_{\rm{200c}}$ by $z=0$. We define the inner and outer extents of galaxy protoclusters using $R_{10}$ and $R_{90}$, which represent the radii, measured relative to the center of mass, that enclose $10\%$ and $90\%$ of the total stellar mass traced by the baseline protocluster population. Lastly, we create density maps using adaptive binning weighted by nearest neighbor distance, which is qualitatively similar to results from a Kernel Density Estimation approach. With this information we investigate how the typical angular sizes used to define spectroscopically-confirmed protocluster samples compare to theoretical expectations, the impact of stellar mass and SFR incompleteness on identifying the highest galaxy density peak in protoclusters, and the relationship between local galaxy density variations in protoclusters and average galaxy properties. Our main conclusions are as follows:
  1. (i)
    The field of view of observational surveys typically used to define spectroscopically-confirmed protoclusters (about $\sim 4\arcmin$) are generally much smaller than the $8\arcmin$ apertures needed to study the progenitors of the most massive galaxy clusters at $3\leq z\leq 5$.
  2. (ii)
    Angular sizes of $4\arcmin$, which are typical in defining spectroscopically-confirmed protocluster, fall within a region of confusion where they could be probing either the innermost region of a massive cluster progenitor or the full extent of a low-mass cluster progenitor.
  3. (iii)
    Stellar mass-limited (${M}_{\star}>10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) and SFR-limited (SFR$>10\leavevmode\nobreak\ {\rm M}_{\odot}\leavevmode\nobreak\ \mathrm{yr^{-1}}$) protocluster populations recover only $\lesssim 40\%$ of the true highest galaxy density peaks within an accuracy of 1 pMpc ($\sim 2-2.5\leavevmode\nobreak\ \arcmin$) at $3\leq z\leq 5$, indicating that the bulk of the densest regions probed by observationally-limited protocluster samples are misidentified.
  4. (iv)
    For protocluster populations complete down to ${M}_{\star}\sim 10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$, inferences about whether the highest galaxy density peak corresponds to a region of accelerated stellar mass growth and enhanced star formation are sensitive to how the density map is weighted. Density maps that are agnostic to stellar mass typically identify highest galaxy density peaks that do not generally correspond to regions of accelerated galaxy evolution, whereas applying stellar mass weighting shifts these peaks toward regions of enhanced star formation and stellar mass growth. This disconnect likely arises because, at $3\leq z\leq 5$, the highest galaxy density peaks do not generally coincide with the highest matter density peaks.
  5. (v)
    Our results suggest that observed protocluster density maps, which are inherently more heavily weighted toward more massive galaxies due to the incompleteness of lower-mass galaxies (e.g., ${M}_{\star}<10^{9.5}\leavevmode\nobreak\ {\rm M}_{\odot}$) at $z>2$, may overestimate the frequency with which the highest density regions are sites of accelerated galaxy evolution relative to the rest of the protocluster.
This investigation focused on the stellar mass and SFR limitations of protocluster populations, but future studies could examine the impact of incompleteness in other observable quantities, such as luminosity and color, which can be modeled with simulations like TNG-Cluster. Additionally, simulation studies that begin with the observer’s definition of protoclusters as high-redshift overdensities — without a priori knowledge of their eventual evolution into clusters — will help elucidate how selection functions and incompleteness influence the fate of high-redshift overdensities identified in observations. Furthermore, detailed comparisons using mock observations, accounting for uncertainties in protocluster membership and selection functions, will be helpful in refining our interpretation of observed protocluster samples and prepare for future surveys (e.g., Euclid Collaboration et al., 2025b). These efforts are timely, as combining the ultra-deep imaging, wide fields of view, and high-resolution spectroscopy from recently launched and next-generation observatories such as JWST, LSST, Euclid, and Roman will lead to an explosion in the number of spectroscopically-confirmed and heterogeneously-selected protocluster samples, significantly enhancing our window into the earliest stages of galaxy cluster formation.
DCB is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303800. DCB is also supported by the UC Chancellor’s Postdoctoral Fellowship. ALC is supported by the Ingrid and Joseph W. Hibben endowed chair at UC San Diego. DN acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG) through an Emmy Noether Research Group (grant number NE 2441/1-1). AP acknowledges funding from the European Union (ERC, COSMIC-KEY, 101087822, PI: Pillepich) The TNG-Cluster simulation has been executed on several machines: with compute time awarded under the TNG-Cluster project on the HoreKa supercomputer, funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research; the bwForCluster Helix supercomputer, supported by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG; the Vera cluster of the Max Planck Institute for Astronomy (MPIA), as well as the Cobra and Raven clusters, all three operated by the Max Planck Computational Data Facility (MPCDF); and the BinAC cluster, supported by the High Performance and Cloud Computing Group at the Zentrum für Datenverarbeitung of the University of Tübingen, the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 37/935-1 FUGG. Many thanks to Maria Polleta for providing feedback that has greatly enhanced the quality of this manuscript. This research made extensive use of Astropy, a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013a, 2018a). Additionally, the Python packages NumPy (Van Der Walt et al., 2011), iPython (Pérez & Granger, 2007), SciPy (Virtanen et al., 2020), and matplotlib (Hunter, 2007) were utilized for our data analysis and presentation. In addition, this research has made use of NASA’s Astrophysics Data System Bibliographic Services.
Appendix A Observational Feasibility of Protocluster Definition Employed in This Work
In Fig. 14, we explore how our definition of simulated protoclusters – galaxies that will reside within the virial radius ($R_{\rm{200c}}$) at $z=0$ – compares with observational scenarios, where such information is inaccessible. Specifically, we compare the contamination fraction, defined as the ratio of interlopers to the sum of protocluster members and interlopers, as a function of separation from the center of mass of the protocluster. Here, interlopers are defined as galaxies that reside within the Lagrangian volume (traced by the protocluster members) but will not reside within $R_{\rm{200c}}$ at $z=0$. In contrast, protocluster members are galaxies that will reside within $R_{\rm{200c}}$ at $z=0$. Both interlopers and protocluster members are constrained to have stellar masses greater than $10^{8.5}\leavevmode\nobreak\ {\rm M}_{\odot}$.
For the six protoclusters shown in Fig. 14 (the same as those in Fig. 2), we find that while our protocluster definition results in some contamination, the level of contamination depends strongly on redshift, separation from the center of mass, and the halo mass of the galaxy cluster at $z=0$. Specifically, contamination is more prevalent at lower redshifts, larger separations, and for progenitors of more massive clusters. At $z>2$, the contamination fraction up to the 90th percentile of member separations is generally less than $20\%$ for the progenitors of the most massive galaxy clusters at $z=0$. These results highlight the existing challenge of establishing a standardized theoretical and observational definition for galaxy protoclusters.
Figure 14: The contamination fraction, defined as the ratio of interlopers to the sum of interlopers and protocluster members, as a function of separation from the center of mass of the protocluster. The gray bands indicate the interquartile range (25th to 75th percentiles) of the center-of-mass-centric separation, while the vertical dashed line shows the median. The 90th percentile is marked by a vertical solid line. These results demonstrate that while the definition used to identify protocluster populations in this work is theoretically sound, it is not observationally perfect, as it would result in some level of contamination from galaxies that will not reside within $R_{\rm{200c}}$ at $z=0$. Empty panels represent cases where the contamination fraction is effectively zero.
Appendix B Table of Spectroscopically-confirmed Protoclusters
Table 1: Spectroscopically-confirmed Protoclusters at $z>2$ with $N_{\mathrm{spec}}>10$
Namea Redshiftb Nspecc Overdensity Tracerd FoV/Aperturee Angular SizeF Referenceg
PHzG237.01+42.50 2.16 31 SMGs/HAEs $10\times 11$ 5.3 Polletta et al. (2021)
PKS1138-262 2.16 54 LAE/HAE/SMG $7\times 7$ 3.5 Pentericci et al. (1998)
BOSS1542 2.24 36 HAE $7\times 7$ 5.3 Shi et al. (2021a)
BOSS1244 2.24 46 HAE $20.4\times 20.4$ 10.2 Shi et al. (2021a)
HS1700FLD 2.30 19 BX/SMG $8\times 8$ 4.0 Lacaille et al. (2019)
BOSS1441 2.32 20 LAE/HAE $6\times 6$ 3 Cai et al. (2017)
USS1558-003 2.53 19 HAE $7\times 4$ 2.8 Pérez-Martínez et al. (2024a)
PCL1002 2.45 11 Spec/LAE/SMG $\pi\times 2.8^{2}$ 2.8 Casey et al. (2015)
4C23.56 2.49 21 HAE $\pi\times 0.4^{2}$ 0.4 Lee et al. (2017)
Surabhi 2.80 17 Spec $39\times 28.5$ 7.6 Shah et al. (2024)
HS1549 2.85 26 LBG/SMG $\pi\times 1.5^{2}$ 1.5 Lacaille et al. (2019)
MRC0052-241 2.86 37 LAE $7\times 7$ 3.5 Venemans et al. (2007)
P2Q1 2.90 12 Spec $7\times 8$ 3.8 Cucciati et al. (2014)
MRC0943-242 2.92 28 LAE $7\times 7$ 3.5 Venemans et al. (2007)
Drishti 2.67 40 Spec $39\times 28.5$ 7.8 Shah et al. (2024)
MRC0316-257 3.13 31 LAE $7\times 7$ 3.5 Venemans et al. (2007)
TNJ2009-3040 3.16 11 LAE $7\times 7$ 3.5 Venemans et al. (2007)
SSA22FLD 3.09 15 LBG/LAE/SMG $11.5\times 9$ 5.1 Steidel et al. (1998)
ClJ0227-0421 3.29 19 Spec $\pi\times 6.2^{2}$ 6.2 Lemaux et al. (2014)
Shrawan 3.30 17 Spec $39\times 28.5$ 9.0 Shah et al. (2024)
MAGAZ3NE J095924+022537 3.36 14 UMG $\pi\times 5^{2}$ 5.0 McConachie et al. (2022)
Smruti 3.47 55 Spec $39\times 28.5$ 8.3 Shah et al. (2024)
D4GD01 3.67 11 LBG $\pi\times 1.8^{2}$ 1.8 Toshikawa et al. (2016)
Sparsh 3.70 22 Spec $39\times 28.5$ 6.2 Shah et al. (2024)
HSC-SSP J100139+022803 3.70 13 LAE $\pi\times 1.8^{2}$ 1.8 Toshikawa et al. (2025b)
PC217.96+32.3 3.79 65 LAE $\pi\times 1.2^{2}$ 1.2 Dey et al. (2016)
DRC-protocluster 4.00 10 DSFG $0.61\times 0.73$ 0.033 Oteo et al. (2018)
TNJ1338-1942 4.11 37 LAE/LBG $7\times 7$ 3.5 Overzier et al. (2008)
Ruchi 4.14 11 Spec $39\times 28.5$ 10 Shah et al. (2024)
SPT2349-56 4.31 14 SMG $\pi\times 0.16^{2}$ 0.16 Miller et al. (2018)
HDF850.1 5.18 23 SMG $7.5\times 6$ 3.4 Calvi et al. (2023)
JADES-GS-OD-5.386 5.4 39 HAE $8.2\times 8.6$ 4.2 Helton et al. (2024b)
z57OD 5.69 44 LAE $\pi\times 4.2^{2}$ 4.2 Harikane et al. (2019)
ODz5p8 5.75 25 Spec $\pi\times 1.42^{2}$ 1.4 Morishita et al. (2024)
PCz6.05-1 5.98 10 LBG $16.7\times 5$ 5.4 Brinch et al. (2024)
SDF 6.01 10 LBG $6\times 6$ 3 Toshikawa et al. (2014)
HSC-z7PCC26 6.54 14 LAE $\pi\times 4.2^{2}$ 4.2 Higuchi et al. (2019)
ClJ1001 6.6 52 Spec $7\times 5$ 3 Champagne et al. (2025)
z66OD 6.59 12 LAE $\pi\times 4.2^{2}$ 4.2 Harikane et al. (2019)
LAGER-z7OD1 6.93 16 LAE $26\times 12$ 9.5 Hu et al. (2021)
a Name of the protocluster as used in the literature.
b Redshift of the protocluster.
c Number of spectroscopically-confirmed protocluster members.
d Galaxy populations used to trace the overdensity: LAE = Lyman-alpha emitter, HAE = narrowband H-alpha emitter, LBG = Lyman break galaxy, BX = ”BX” galaxy of Adelberger et al. (2005), SMG = submillimeter galaxy, UMG = ultramassive galaxy, Spec = spectroscopic survey.
e Field of view or aperture used to identify the protocluster, in units of arcminutes2.
f Angular radius of the protocluster, defined as half of the average of the field of view’s length and width, or the radius of the aperture, in arcminutes.
g References for studies where the protocluster was first reported or recently analyzed, including those expanding its spectroscopic membership beyond ten.